Base on the stoichiometry road map calculate the problem below by showing the step by step solution.

Given the following reaction:

Al2 (SO4)3 + 6 NaOH --------> 2 Al(OH)3 + 3 Na2SO4

What mass of Al(OH)3 is produced if 22.7 g of NaOH is consumed?


1
Expert's answer
2021-10-17T11:21:27-0400

v(NaOH)=22.7/40=0,5675mol

v(Al(OH)4)=0.5675/3=0.19mol

m(Al(OH)3)=0.19*78=14.82g